OpenStudy (anonymous):

Elimination Method. 6x-2y=12 -5x+9=56

OpenStudy (anonymous):

Showing work that's what I got but I'm not sure. Can somebody please correct me? 5(6x-2y)=5(12) 6(-5x+9)=6(56) 30x-10y=60 -30x+54=336

OpenStudy (nincompoop):

good

OpenStudy (nincompoop):

start "eliminating"

OpenStudy (anonymous):

Help. ? Lol

OpenStudy (anonymous):

Solve for the y, you can see you have 30x and -30x, so it will get cancelled out and now you can solve for y and afterwards you can pick any of the following equations and plug the y in to the get the corresponding x value.

OpenStudy (hockeychick23):

ok so we start with: 5(6x-2y)=5(12) 6(-5x+9)=6(56) Now lets distribute the 5 in the first equation and 6 in the second equation. 30x-10y=60 -30x+54= 336 So we already have a variable we can eliminate here. When we add the second equation to the first we get: 30x+ (-30x) - 10y+54= 336+60 Do you know what to do next?